# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 4.1 Action in Fourier space
$\pi^{M}(x)=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}e^{-i\omega t+i\mathbf{k}\cdot\mathbf{x}}\pi^{M}_{\omega,\mathbf{k}},$ (180)
with $M=\mu,4$ and we assigned $M=4$ for $\theta$. In the plane wave mode analysis, ${\rm Im}~{}\omega>$ implies the exponential growth of perturbations; therefore, the criteria of stability is given by ${\rm Im}~{}\omega\leq 0$. Hereafter, we consider Fourier space, and the arguments for the NG modes are suppressed for notational simplicity. Then, the dynamical action is recast into
$I^{(1,0,1)}_{d}+I^{(1,1,1)}_{d}$ $=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}({\bar{\pi}}_{0}^{*}~{}{\bar{\pi}}_{i}^{*}~{}\bar{\theta}^{*})2\mathcal{E}\begin{pmatrix}\pi^{0}\\ \pi^{j}\\ \theta\end{pmatrix},$ (181)
@@
$+ik^{2}\omega\left(\beta^{2}\mu^{2}\chi-\beta\mu(\tau_{1}+\tau_{3})+\kappa\right),$ (182)
$2\mathcal{E}^{0}{}_{i}$ $=k_{i}\omega(\rho+P-\mu n)+ik_{i}\omega^{2}(\kappa_{2}+\lambda_{1}-\beta\mu(\nu_{2}+\tau_{2})),$ (183)
$2\mathcal{E}^{0}{}_{4}$ $=\omega^{2}(\mu n_{\mu}-\rho_{\mu})+i\beta\omega^{3}\left(\beta\mu\nu_{3}-\nu_{4}\right)+i\beta k^{2}\omega\left(\beta\mu\chi_{1}-\tau_{3}\right),$ (184)
$2\mathcal{E}^{i}{}_{0}$ $=\beta\omega k_{i}P_{\beta}-ik_{i}\omega^{2}(\kappa_{1}+\lambda_{2}-\beta\mu(\nu_{5}+\tau_{3})),$ (185)
$2\mathcal{E}^{i}{}_{j}$ $=-\left(\omega^{2}(\rho+\rho)+i\kappa\omega^{3}+i\eta k^{2}\omega\right)\delta^{i}{}_{j}-\frac{1}{3}ik_{i}k_{j}\omega(3\zeta+\eta),,$ (186)
$2\mathcal{E}^{i}{}_{4}$ $=k_{i}P_{\mu}\omega+i\beta k_{i}\omega^{2}(\nu_{5}+\tau_{3}),$ (187)
$2\mathcal{E}^{4}{}_{0}$ $=i\omega^{3}\left(\beta^{2}\mu\nu_{3}-\beta\nu_{1}\right)+i\beta\omega k^{2}\left(\beta\mu\chi_{1}-\tau_{1}\right)+\beta n_{\beta}\omega^{2},$ (188)
$2\mathcal{E}^{4}{}_{i}$ $=-\omega k_{i}n-i\beta\omega^{2}k_{i}(\nu_{2}+\tau_{2}),$ (189)
$2\mathcal{E}^{4}{}_{4}$ $=n_{\mu}\omega^{2}+i\beta^{2}\omega^{3}\nu_{3}+i\beta^{2}\omega k^{2}\chi_{1}.$ (190)
We are interested in thermal states, so the KMS condition is assumed. However, we explicitly distinguish the equivalent hydrodynamical coefficients, such as $\lambda_{1}$ and $\lambda_{2}$, to discuss field redefinition later. The thermodynamical relations (272) and integrability condition (273) yield
$\mathcal{E}^{0}{}_{i}=-\mathcal{E}^{i}{}_{0},~{}\mathcal{E}^{0}{}_{4}=\mathcal{E}^{4}{}_{0},~{}\mathcal{E}^{i}{}_{4}=-\mathcal{E}^{4}{}_{i}.$ (191)
@@
$I^{(2,0,0)}_{d}$ $=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}({\bar{\pi}}_{0}^{*}~{}{\bar{\pi}}_{i}^{*}~{}{\bar{\theta}}^{*})\mathcal{N}\begin{pmatrix}\bar{\pi}_{0}\\ \bar{\pi}_{j}\\ \bar{\theta}\end{pmatrix},$ (192)
@@
$\mathcal{N}^{00}$ $=\frac{i}{\beta}\left(\omega^{2}(\mu^{2}\nu_{3}-2\mu\nu_{1}+\epsilon)+(\kappa+\mu^{2}\chi-2\mu\tau)k^{2}\right),$ (193)
$\mathcal{N}^{0i}$ $=-\frac{i}{\beta}\omega k^{i}(\lambda+\kappa-\mu\tau-\mu\nu_{2}),$ (194)
$\mathcal{N}^{04}$ $=\frac{i}{\beta}(k^{2}(-\tau+\mu\chi)+\omega^{2}(-\nu_{1}+\mu\nu_{3})),$ (195)
$\mathcal{N}^{ij}$ $=\frac{i}{\beta}\left[(\kappa\omega^{2}+\eta k^{2})\delta^{ij}+\left(\zeta+\frac{1}{3}\eta\right)k^{i}k^{j}\right],$ (196)
$\mathcal{N}^{i4}$ $=\frac{i}{\beta}\omega k^{i}\left[\tau+\nu_{2}\right],$ (197)
$\mathcal{N}^{44}$ $=\frac{i}{\beta}\left[\omega^{2}\nu_{3}+k^{2}\chi\right].$ (198)
@@ 4.2 Constraint equations
Combining the dynamical and noise actions, we find
$I_{d}=$ $\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}\bm{\pi}^{\dagger}G^{-1}\bm{\pi},$ (199)
@@
$\bm{\pi}=\begin{pmatrix}\pi^{\mu}\\ \theta\\ \bar{\pi}_{\mu}\\ \bar{\theta}\end{pmatrix},~{}G^{-1}=\begin{pmatrix}0&\mathcal{E}^{*}\\ \mathcal{E}&\mathcal{N}\end{pmatrix}.$ (200)
@@
$\frac{\delta I_{d}}{\delta\pi^{0*}}$ $={\bar{\pi}}_{0}(\mathcal{E}^{0}{}_{0})^{*}+{\bar{\pi}}_{i}(\mathcal{E}^{i}{}_{0})^{*}+\bar{\theta}(\mathcal{E}^{4}{}_{0})^{*}=0,$ (201)
$\frac{\delta I_{d}}{\delta\theta^{*}}$ $={\bar{\pi}}_{0}(\mathcal{E}^{0}{}_{4})^{*}+{\bar{\pi}}_{i}(\mathcal{E}^{i}{}_{4})^{*}+\bar{\theta}(\mathcal{E}^{4}{}_{4})^{*}=0,$ (202)
$\frac{\delta I_{d}}{\delta\bar{\pi}^{*}_{0}}$ $=2\mathcal{E}^{0}{}_{0}\pi^{0}+2\mathcal{E}^{0}{}_{i}\pi^{i}+2\mathcal{E}^{0}{}_{4}\theta+\mathcal{N}^{00}{\bar{\pi}}_{0}+\mathcal{N}^{0i}{\bar{\pi}}_{i}+\mathcal{N}^{04}{\bar{\theta}}=0,$ (203)
$\frac{\delta I_{d}}{\delta\bar{\theta}^{*}}$ $=2\mathcal{E}^{4}{}_{0}\pi^{0}+2\mathcal{E}^{4}{}_{i}\pi^{i}+2\mathcal{E}^{4}{}_{4}\theta+\mathcal{N}^{40}{\bar{\pi}}_{0}+\mathcal{N}^{4i}{\bar{\pi}}_{i}+\mathcal{N}^{44}{\bar{\theta}}=0,$ (204)
where, as defined above, the 4th index corresponds to $\theta$. Note that the complex conjugates are not regarded as independent variables when taking the variation since the NG modes are real in real space. For classical dynamics, the noise field is set to 0 so that the first two equations are trivial. Then the third and fourth equations reduce to the linearized continuity equation and number conservation. In the present case, we want to eliminate the redundant variables up to the noise order consistently, so we need to account for the noise correction in Eqs. (203) and (204). These equations are solved as
$\pi^{0}$ $=L^{0}{}_{i}\pi^{i}+M^{0i}\bar{\pi}_{i},$ (205)
$\theta$ $=L^{4}{}_{i}\pi^{i}+M^{4i}\bar{\pi}_{i},$ (206)
$\bar{\pi}_{0}$ $=\bar{\pi}_{i}L^{i}{}_{0},$ (207)
$\bar{\theta}$ $=\bar{\pi}_{i}L^{i}{}_{4},$ (208)
@@
$L^{0}{}_{i}$ $=-\frac{\mathcal{E}^{0}{}_{i}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{i}}{\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{0}},$ (209)
$L^{4}{}_{i}$ $=\frac{\mathcal{E}^{0}{}_{i}\mathcal{E}^{4}{}_{0}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{i}}{\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{0}},$ (210)
$L^{i}{}_{0}$ $=-(L^{0}{}_{i})^{*},$ (211)
$L^{i}{}_{4}$ $=-(L^{4}{}_{i})^{*},$ (212)
@@
$+\frac{\mathcal{N}^{04}(-(\mathcal{E}^{0}{}_{0})^{*}\mathcal{E}^{4}{}_{4}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{4}{}_{i})^{*}+\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{i})^{*}(\mathcal{E}^{4}{}_{4})^{*}+(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{0}{}_{i})^{*}\mathcal{E}^{4}{}_{4})}{\left((\mathcal{E}^{0}{}_{4})^{2}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}\right)\left((\mathcal{E}^{0}{}_{4})^{*2}-(\mathcal{E}^{0}{}_{0})^{*}(\mathcal{E}^{4}{}_{4})^{*}\right)},$ (213)
@@
$+\frac{\mathcal{N}^{04}(\mathcal{E}^{0}{}_{0}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{0}(\mathcal{E}^{0}{}_{i})^{*}(\mathcal{E}^{4}{}_{4})^{*}+(\mathcal{E}^{0}{}_{0})^{*}\mathcal{E}^{0}{}_{4}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{0}{}_{i})^{*})}{\left((\mathcal{E}^{0}{}_{4})^{2}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}\right)\left((\mathcal{E}^{0}{}_{4})^{*2}-(\mathcal{E}^{0}{}_{0})^{*}(\mathcal{E}^{4}{}_{4})^{*}\right)}.$ (214)
@@
$=$ $2{\bar{\pi}}_{i}^{*}\tilde{\mathcal{E}}^{i}{}_{j}\pi^{j}+{\bar{\pi}}_{i}^{*}\tilde{\mathcal{N}}^{ij}\bar{\pi}_{j},$ (215)
@@
~$\tilde{\mathcal{E}}^{i}{}_{j}$ $=\mathcal{E}^{i}{}_{j}+(L^{i}{}_{M})^{*}\mathcal{E}^{M}{}_{j},$ (216)
$\tilde{\mathcal{N}}^{ij}$ $=\mathcal{N}^{ij}+\mathcal{N}^{iN}L^{j}{}_{N}+(L^{i}{}_{M})^{*}\mathcal{N}^{Mj}+(L^{i}{}_{M})^{*}\mathcal{N}^{MN}L^{j}{}_{N}.$ (217)
The effective action is further simplified by separating the longitudinal and transverse modes. Consider the scalar-vector decomposition in flat space Kodama:1984ziu
${\bar{\pi}}_{i}=ik_{i}{\bar{\pi}}^{\rm L}+\bar{\pi}^{\rm T}_{i},~{}ik^{i}\bar{\pi}^{\rm T}_{i}=0.$ (218)
Since the noise NG modes are linear representations of spacetime diffeomorphism, the scalar vector decomposition in flat space is straightforward. However, we should note that the index of the physical NG mode is a label of four scalar fields, and hence, it cannot be regarded as a spacetime index of a 4-vector. In fact, the diffeomorphism transformation for the NG modes is nonlinear:
$\pi^{\prime\mu}(x^{\prime})=\pi^{\mu}(x)+\xi^{\mu}(x).$ (219)
The full effective action is covariant under this transformation. Nevertheless, one may also consider the scalar-vector decomposition in Eq. (219) for the physical NG mode for the following reasons. When expanding the action in terms of the NG modes, as done in Sec. 3.2, $\pi^{i}$ always appears contracted with 3-vectors. For example, $I_{d}^{(1,0,1)}\supset-\dot{\bar{\pi}}_{i}\dot{\pi}^{i}(\rho+P)$. Diffeomorphism invariance is nonlinearly realized by combining all terms in the full effective action. However, order by order in the NG mode expansion, each term can be regarded as a covariant block in flat space. Therefore, one may assign a linear transformation to $\pi^{i}$, allowing the spatial index to be treated as if it belongs to a linear representation of spatial diffeomorphism.
@@
$I_{\pi}=$ $\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}\tilde{\bm{\pi}}^{\dagger}\tilde{G}^{-1}\tilde{\bm{\pi}},$ (220)
@@
$\tilde{G}^{-1}=\begin{pmatrix}0&0&\tilde{\mathcal{E}}^{{\rm T}*}&0&\\ 0&0&0&\tilde{\mathcal{E}}^{{\rm L}*}\\ \tilde{\mathcal{E}}^{{\rm T}}&0&\tilde{\mathcal{N}}^{\rm T}&0\\ 0&\tilde{\mathcal{E}}^{{\rm L}}&0&\tilde{\mathcal{N}}^{\rm L}\end{pmatrix}.$ (222)

### v3

@@ 5.1 Action in Fourier space
$\pi^{M}(x)=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}e^{-i\omega t+i\mathbf{k}\cdot\mathbf{x}}\pi^{M}_{\omega,\mathbf{k}},$ (5.1)
with $M=\mu,4$ and we assigned $M=4$ for $\theta$. In the plane wave mode analysis, ${\rm Im}~\omega>0$ implies the exponential growth of perturbations; therefore, the criteria of stability is given by ${\rm Im}~\omega\leq 0$. Hereafter, we consider Fourier space, and the arguments for the NG modes are suppressed for notational simplicity. Then, the dynamical action is recast into
$I^{(1,0,1)}_{d}+I^{(1,1,1)}_{d}$ $=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}({\bar{\pi}}_{0}^{*}~{\bar{\pi}}_{i}^{*}~\bar{\theta}^{*})2\mathcal{E}\begin{pmatrix}\pi^{0}\\ \pi^{j}\\ \theta\end{pmatrix},$ (5.2)
@@
$+ik^{2}\omega\left(\beta^{2}\mu^{2}\chi-\beta\mu(\tau_{1}+\tau_{3})+\kappa\right),$ (5.3)
$2\mathcal{E}^{0}{}_{i}$ $=k_{i}\omega(\rho+P-\mu n)+ik_{i}\omega^{2}(\kappa_{2}+\lambda_{1}-\beta\mu(\nu_{2}+\tau_{2})),$ (5.4)
$2\mathcal{E}^{0}{}_{4}$ $=\omega^{2}(\mu n_{\mu}-\rho_{\mu})+i\beta\omega^{3}\left(\beta\mu\nu_{3}-\nu_{4}\right)+i\beta k^{2}\omega\left(\beta\mu\chi_{1}-\tau_{3}\right),$ (5.5)
$2\mathcal{E}^{i}{}_{0}$ $=\beta\omega k_{i}P_{\beta}-ik_{i}\omega^{2}(\kappa_{1}+\lambda_{2}-\beta\mu(\nu_{5}+\tau_{3})),$ (5.6)
$2\mathcal{E}^{i}{}_{j}$ $=-\left(\omega^{2}(\rho+\rho)+i\kappa\omega^{3}+i\eta k^{2}\omega\right)\delta^{i}{}_{j}-\frac{1}{3}ik_{i}k_{j}\omega(3\zeta+\eta),,$ (5.7)
$2\mathcal{E}^{i}{}_{4}$ $=k_{i}P_{\mu}\omega+i\beta k_{i}\omega^{2}(\nu_{5}+\tau_{3}),$ (5.8)
$2\mathcal{E}^{4}{}_{0}$ $=i\omega^{3}\left(\beta^{2}\mu\nu_{3}-\beta\nu_{1}\right)+i\beta\omega k^{2}\left(\beta\mu\chi_{1}-\tau_{1}\right)+\beta n_{\beta}\omega^{2},$ (5.9)
$2\mathcal{E}^{4}{}_{i}$ $=-\omega k_{i}n-i\beta\omega^{2}k_{i}(\nu_{2}+\tau_{2}),$ (5.10)
$2\mathcal{E}^{4}{}_{4}$ $=n_{\mu}\omega^{2}+i\beta^{2}\omega^{3}\nu_{3}+i\beta^{2}\omega k^{2}\chi_{1}.$ (5.11)
We are interested in thermal states, so the KMS condition is assumed. However, we explicitly distinguish the equivalent hydrodynamical coefficients, such as $\lambda_{1}$ and $\lambda_{2}$, to discuss field redefinition later. The thermodynamical relations (B.2) and integrability condition (B.3) yield
$\mathcal{E}^{0}{}_{i}=-\mathcal{E}^{i}{}_{0},~\mathcal{E}^{0}{}_{4}=\mathcal{E}^{4}{}_{0},~\mathcal{E}^{i}{}_{4}=-\mathcal{E}^{4}{}_{i}.$ (5.12)
@@
$I^{(2,0,0)}_{d}$ $=\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}({\bar{\pi}}_{0}^{*}~{\bar{\pi}}_{i}^{*}~{\bar{\theta}}^{*})\mathcal{N}\begin{pmatrix}\bar{\pi}_{0}\\ \bar{\pi}_{j}\\ \bar{\theta}\end{pmatrix},$ (5.13)
@@
$\mathcal{N}^{00}$ $=\frac{i}{\beta}\left(\omega^{2}(\mu^{2}\nu_{3}-2\mu\nu_{1}+\epsilon)+(\kappa+\mu^{2}\chi-2\mu\tau)k^{2}\right),$ (5.14)
$\mathcal{N}^{0i}$ $=-\frac{i}{\beta}\omega k^{i}(\lambda+\kappa-\mu\tau-\mu\nu_{2}),$ (5.15)
$\mathcal{N}^{04}$ $=\frac{i}{\beta}(k^{2}(-\tau+\mu\chi)+\omega^{2}(-\nu_{1}+\mu\nu_{3})),$ (5.16)
$\mathcal{N}^{ij}$ $=\frac{i}{\beta}\left[(\kappa\omega^{2}+\eta k^{2})\delta^{ij}+\left(\zeta+\frac{1}{3}\eta\right)k^{i}k^{j}\right],$ (5.17)
$\mathcal{N}^{i4}$ $=\frac{i}{\beta}\omega k^{i}\left[\tau+\nu_{2}\right],$ (5.18)
$\mathcal{N}^{44}$ $=\frac{i}{\beta}\left[\omega^{2}\nu_{3}+k^{2}\chi\right].$ (5.19)
Our effective action is found to the noise order, which is a new result of this work.
@@ 5.2 Constraint equations
In the traditional analysis, one often computes the determinant of $\mathcal{E}$, which is recast into
$\det(\mathcal{E})\propto\det(\mathcal{E}^{\rm long.})\det(\mathcal{E}^{\rm trans.}),$ (5.20)
where the right hand side implies the product of the determinants of the longitudinal and transverse modes. The dispersion relations are found from $\det(\mathcal{E})=0$. The redundancy in the variables suggests the necessity of solving the constraint equations. In this section, we carefully solve the constraint equations given by the continuity equation and change conservation.
In the action formalism, the dynamical and noise actions are written as
$I_{d}=$ $\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}\bm{\pi}^{\dagger}G^{-1}\bm{\pi},$ (5.21)
@@
$\bm{\pi}=\begin{pmatrix}\pi^{\mu}\\ \theta\\ \bar{\pi}_{\mu}\\ \bar{\theta}\end{pmatrix},~G^{-1}=\begin{pmatrix}0&\mathcal{E}^{*}\\ \mathcal{E}&\mathcal{N}\end{pmatrix}.$ (5.22)
@@
$\frac{\delta I_{d}}{\delta\pi^{0*}}$ $={\bar{\pi}}_{0}(\mathcal{E}^{0}{}_{0})^{*}+{\bar{\pi}}_{i}(\mathcal{E}^{i}{}_{0})^{*}+\bar{\theta}(\mathcal{E}^{4}{}_{0})^{*}=0,$ (5.23)
$\frac{\delta I_{d}}{\delta\theta^{*}}$ $={\bar{\pi}}_{0}(\mathcal{E}^{0}{}_{4})^{*}+{\bar{\pi}}_{i}(\mathcal{E}^{i}{}_{4})^{*}+\bar{\theta}(\mathcal{E}^{4}{}_{4})^{*}=0,$ (5.24)
$\frac{\delta I_{d}}{\delta\bar{\pi}^{*}_{0}}$ $=2\mathcal{E}^{0}{}_{0}\pi^{0}+2\mathcal{E}^{0}{}_{i}\pi^{i}+2\mathcal{E}^{0}{}_{4}\theta+\mathcal{N}^{00}{\bar{\pi}}_{0}+\mathcal{N}^{0i}{\bar{\pi}}_{i}+\mathcal{N}^{04}{\bar{\theta}}=0,$ (5.25)
$\frac{\delta I_{d}}{\delta\bar{\theta}^{*}}$ $=2\mathcal{E}^{4}{}_{0}\pi^{0}+2\mathcal{E}^{4}{}_{i}\pi^{i}+2\mathcal{E}^{4}{}_{4}\theta+\mathcal{N}^{40}{\bar{\pi}}_{0}+\mathcal{N}^{4i}{\bar{\pi}}_{i}+\mathcal{N}^{44}{\bar{\theta}}=0,$ (5.26)
where, as defined above, the 4th index corresponds to $\theta$. Note that the complex conjugates are not regarded as independent variables. For classical dynamics, the noise field is set to 0 so that the first two equations are trivial. Then the third and fourth equations reduce to the linearized continuity equation and number conservation. In the present case, we want to eliminate the redundant variables up to the noise order consistently, so we need to account for the noise correction in Eqs. (5.25) and (5.26). These equations are solved as
$\pi^{0}$ $=L^{0}{}_{i}\pi^{i}+M^{0i}\bar{\pi}_{i},$ (5.27)
$\theta$ $=L^{4}{}_{i}\pi^{i}+M^{4i}\bar{\pi}_{i},$ (5.28)
$\bar{\pi}_{0}$ $=\bar{\pi}_{i}L^{i}{}_{0},$ (5.29)
$\bar{\theta}$ $=\bar{\pi}_{i}L^{i}{}_{4},$ (5.30)
@@
$L^{0}{}_{i}$ $=-\frac{\mathcal{E}^{0}{}_{i}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{i}}{\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{0}},$ (5.31)
$L^{4}{}_{i}$ $=\frac{\mathcal{E}^{0}{}_{i}\mathcal{E}^{4}{}_{0}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{i}}{\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}-\mathcal{E}^{0}{}_{4}\mathcal{E}^{4}{}_{0}},$ (5.32)
$L^{i}{}_{0}$ $=-(L^{0}{}_{i})^{*},$ (5.33)
$L^{i}{}_{4}$ $=-(L^{4}{}_{i})^{*},$ (5.34)
@@
$+\frac{\mathcal{N}^{04}(-(\mathcal{E}^{0}{}_{0})^{*}\mathcal{E}^{4}{}_{4}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{4}{}_{i})^{*}+\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{i})^{*}(\mathcal{E}^{4}{}_{4})^{*}+(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{0}{}_{i})^{*}\mathcal{E}^{4}{}_{4})}{\left((\mathcal{E}^{0}{}_{4})^{2}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}\right)\left((\mathcal{E}^{0}{}_{4})^{*2}-(\mathcal{E}^{0}{}_{0})^{*}(\mathcal{E}^{4}{}_{4})^{*}\right)},$ (5.35)
@@
$+\frac{\mathcal{N}^{04}(\mathcal{E}^{0}{}_{0}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{0}(\mathcal{E}^{0}{}_{i})^{*}(\mathcal{E}^{4}{}_{4})^{*}+(\mathcal{E}^{0}{}_{0})^{*}\mathcal{E}^{0}{}_{4}(\mathcal{E}^{4}{}_{i})^{*}-\mathcal{E}^{0}{}_{4}(\mathcal{E}^{0}{}_{4})^{*}(\mathcal{E}^{0}{}_{i})^{*})}{\left((\mathcal{E}^{0}{}_{4})^{2}-\mathcal{E}^{0}{}_{0}\mathcal{E}^{4}{}_{4}\right)\left((\mathcal{E}^{0}{}_{4})^{*2}-(\mathcal{E}^{0}{}_{0})^{*}(\mathcal{E}^{4}{}_{4})^{*}\right)}.$ (5.36)
@@
$=$ $2{\bar{\pi}}_{i}^{*}\tilde{\mathcal{E}}^{i}{}_{j}\pi^{j}+{\bar{\pi}}_{i}^{*}\tilde{\mathcal{N}}^{ij}\bar{\pi}_{j},$ (5.37)
@@
~$\tilde{\mathcal{E}}^{i}{}_{j}$ $=\mathcal{E}^{i}{}_{j}+(L^{i}{}_{M})^{*}\mathcal{E}^{M}{}_{j},$ (5.38)
$\tilde{\mathcal{N}}^{ij}$ $=\mathcal{N}^{ij}+\mathcal{N}^{iN}L^{j}{}_{N}+(L^{i}{}_{M})^{*}\mathcal{N}^{Mj}+(L^{i}{}_{M})^{*}\mathcal{N}^{MN}L^{j}{}_{N}.$ (5.39)
The effective action is further simplified by separating the longitudinal and transverse modes. Consider the scalar–vector decomposition in flat space [27]
${\bar{\pi}}_{i}=ik_{i}{\bar{\pi}}^{\rm L}+\bar{\pi}^{\rm T}_{i},~ik^{i}\bar{\pi}^{\rm T}_{i}=0.$ (5.40)
However, it is important to note that the rotation group is doubled on the CTP, and the physical rotation group acting on $\bar{\pi}_{i}$ differs from that of a physical 3-vector. As discussed in the previous sections, the physical rotation acting on $\bar{g}_{\mu\nu}$ is trivial, and thus $\bar{\pi}_{i}$ transforms as a scalar under spatial rotations. In contrast, the physical rotation acting on $\pi^{i}$ is nonlinearly realized and corresponds to a constant shift. Indeed, products involving $\pi^{i}$ and $\bar{\pi}_{i}$ appear in expressions such as $I_{d}^{(1,0,1)}\supset-\dot{\bar{\pi}}_{i}\dot{\pi}^{i}(\rho+P)$. The shift symmetry in $\pi^{i}$ ensures that such building blocks are invariant under spatial rotations. Therefore, it is consistent to assign a linear transformation property to $\pi^{i}$ and $\bar{\pi}_{i}$, allowing its spatial index to be treated as if it belongs to a linear representation of the rotation group.
@@
$I_{\pi}=$ $\int\frac{d\omega}{2\pi}\int\frac{d^{3}k}{(2\pi)^{3}}\tilde{\bm{\pi}}^{\dagger}\tilde{G}^{-1}\tilde{\bm{\pi}},$ (5.41)
@@
$\tilde{G}^{-1}=\begin{pmatrix}0&0&\tilde{\mathcal{E}}^{{\rm T}*}&0&\\ 0&0&0&\tilde{\mathcal{E}}^{{\rm L}*}\\ \tilde{\mathcal{E}}^{{\rm T}}&0&\tilde{\mathcal{N}}^{\rm T}&0\\ 0&\tilde{\mathcal{E}}^{{\rm L}}&0&\tilde{\mathcal{N}}^{\rm L}\end{pmatrix}.$ (5.43)